OpenStudy (anonymous):

Show sin^4theta-cos^4theta=sin^2theta-cos^2theta

OpenStudy (amorfide):

you know that \[\sin^{4}\theta = (\sin^{2}\theta)^{2}\] \[\cos^{4}\theta = (\cos^{2}\theta)^{2}\] you know \[\sin^{2}\theta=1-\cos^{2}\theta\] \[\cos^{2}\theta = 1-\sin^{2}\theta\] by using the identity \[\sin^{2}\theta+\cos^{2}\theta=1\] it is just a matter of substituting in what you know and simplifiying

OpenStudy (anonymous):

thank you!